Find the equilibrium pH and concentrations of H2A, HA and A a solution containing 0.0365 M Na2SO3. n pH [HA] [HA] [A²1
Answer #1

0.0365M NOSO Ab = kw a losunat Ka, 250g = .1 4x102 kb2 = k do 7.14x183 Ka tha Jog = 6,5x108 Na₂Salz - anat tsao ? Soon +H₂O →CHSO4Сон .. Кь, - C4547 7-14 x6/3, 2 sx = -1 (2-5хіб - .: сна-0,1 = = 14 хон сон] = 7.5xjp“. - 7-stoyuk - 7.5xoy ДА 21 - о